# Separation and collapse of equilibria inequalities on AND–OR trees without shape constraints

### v1

@@ Abstract
Herein, we investigate the randomized complexity, which is the least cost against the worst input, of AND–OR tree computation by imposing various restrictions on the algorithm to find the Boolean value of the root of that tree and no restrictions on the tree shape. When a tree satisfies a certain condition regarding its symmetry, directional algorithms proposed by Saks and Wigderson (1986), special randomized algorithms, are known to achieve the randomized complexity. Furthermore, there is a known example of a tree that is so unbalanced that no directional algorithm achieves the randomized complexity (Vereshchagin 1998). In this study, we aim to identify where deviations arise between the general randomized Boolean decision tree and its special case, directional algorithms. In this paper, we show that for any AND–OR tree, randomized depth-first algorithms, which form a broader class compared with directional algorithms, have the same equilibrium as that of the directional algorithms. Thus, we get the collapse result on equilibria inequalities that holds for an arbitrary AND–OR tree. This implies that there exists a case where even depth-first algorithms cannot be the fastest, leading to the separation result on equilibria inequality. Additionally, a new algorithm is introduced as a key concept for proof of the separation result.
@@
2000 MSC: Primary 68T20, 68Q17; Secondary 03D15, 91A60
@@ 1.1 Background
AND–OR tree is a computation model of a read-once Boolean function, that is, a function in which each of its variable appears exactly once in its formula. We are interested in the decision tree complexity of a tree here, i.e., the computational cost is measured by the number of leaves the algorithm (i.e., the decision tree) has queried. $\alpha$ runs the algorithms computing the tree and $\omega$ runs the assignments to the tree, then the decision tree complexity is defined as follows:
@@
We are interested in the case where we allow randomization on algorithms and/or assignments. The expected value of cost is then simply called cost. For a fixed tree $T$, we consider two types of equilibrium. These equilibria are good indicators of the complexity of the problem of finding the value of the root. One type of the equilibrium is the randomized complexity, where $A$ runs randomized algorithms and $\omega$ runs (nonrandomized) assignments to the tree, and is defined as
@@
[6, Theorem 2.1] (see also [11, Lemma 5.1]) Suppose an AND–OR tree $T$ is weakly balanced. Then $R(T)=d(T)$.
@@ 1.2 Our goal: Collapse and separation
Considering Example 1.3, it is natural to ask where does the gap lies between $R$ and $d$. We show a collapse on equilibria $R_{\mathrm{DF}}(T)=d(T)$ for all AND–OR tree $T$: The randomized complexity with respect to the depth-first algorithms equals that with respect to the directional algorithms (RDAs). Here, no shape constraints are imposed on the tree $T$. Any node of $T$ can have any number (${\geq}2$) of child nodes, and there can be a mixture of AND and OR nodes among the child nodes. Thus, we obtain a separation result on the equilibria $R(T)<R_{\mathrm{DF}}(T)$ for a certain $T$: The gap lies between the general algorithms and the depth-first algorithms. In other words, our result implies that there exists a case where a depth-first algorithm cannot attain the randomized complexity $R$. This study thus provides a deeper understanding of the depth-first algorithm for AND–OR trees.
@@ 2 Definitions and Preliminaries
In this section, we will define some basic concepts and see some fundamental facts on AND–OR trees. See also [2, Chapter 12] and [12] for helpful examples.
@@ 3.1 Replacement of sub-algorithms
[11, Lemma 3.1] There exists an RDA on $T$ that attains both $d_{0}(T)$ and $d_{1}(T)$.

### v2

@@ Abstract
Herein, we investigate the zero-error randomized complexity, which is the least cost against the worst input, of AND–OR tree computation by imposing various restrictions on the algorithm to find the Boolean value of the root of that tree and no restrictions on the tree shape. When a tree satisfies a certain condition regarding its symmetry, directional algorithms proposed by Saks and Wigderson (1986), special randomized algorithms, are known to achieve the randomized complexity. Furthermore, there is a known example of a tree that is so unbalanced that no directional algorithm achieves the randomized complexity (Vereshchagin 1998). In this study, we aim to identify where deviations arise between the general randomized Boolean decision tree and its special case, directional algorithms. In this paper, we show that for any AND–OR tree, randomized depth-first algorithms, which form a broader class compared with directional algorithms, have the same equilibrium as that of the directional algorithms. Thus, we get the collapse result on equilibria inequalities that holds for an arbitrary AND–OR tree. This implies that there exists a case where even depth-first algorithms cannot be the fastest, leading to the separation result on equilibria inequality. Additionally, a new algorithm is introduced as a key concept for proof of the separation result.
@@
2000 MSC: Primary 68T20, 68Q17; Secondary 03D15, 91A60;
CCS Concepts: Theory of computation $\to$ Probabilistic computation; Oracles and decision trees;
@@ 1.1 Background
AND–OR tree is a computation model of a read-once Boolean function, that is, a function in which each of its variable appears exactly once in its formula. We are interested in the decision tree complexity (or deterministic complexity) of a tree here, i.e., the computational cost is measured by the number of leaves the algorithm (i.e., the decision tree) has queried. $\alpha$ runs the algorithms computing the tree and $\omega$ runs the assignments to the tree, then the decision tree complexity is defined as follows:
@@
We are interested in the case where we allow randomization on algorithms and/or assignments. The expected value of cost is then simply called cost. For a fixed tree $T$, we consider two types of equilibrium. These equilibria are good indicators of the complexity of the problem of finding the value of the root. One type of the equilibrium is the ((((zero-error)))) randomized complexity, where $A$ runs randomized algorithms and $\omega$ runs (nonrandomized) assignments to the tree, and is defined as
@@
It is well known that the randomized complexity may be smaller than the deterministic complexity [14]. In other words, randomization may improve efficiency. There is a considerable amount of prior research on separation between deterministic complexity and randomized complexity. In particular, an open problem in [14] has been solved in mid 2010s, and remarkable developments continue to take place [2, 10, 9].
Note the differences in symbols in the literature. Our definition of $R$ is the same as [14]. We later define $R_{0}$ and $R_{1}$ also in the style of [14]; let us call this notation “classical style”. On the other hand, in some recent literature [2, 10, 9], the symbols $R$, $R_{0}$, and $R_{1}$ are used to mean something different from what they do in [14]; let us call it “modern style”. Table 1 provides summary of differences. In the row of “zero error, the root is $i$”, the truth assignment is restricted to those with root value $i$, and the algorithm is zero-error regardless of the root value. This paper adopt classical style.
meaning classical style (this paper) modern style
zero error version $R$ $R_{0}$
zeo error, the root is 00 $R_{0}$
zero error, the root is $1$ $R_{1}$
bounded error version $R$
one-sided error version $R_{1}$
Table 1: Symbols for variants of randomized complexity
@@
[7, Theorem 2.1] (see also [14, Lemma 5.1]) Suppose an AND–OR tree $T$ is weakly balanced. Then $R(T)=d(T)$.
@@ 1.2 Our goal: Collapse and separation
Our main motivations for focusing on the directional algorithms is precisely because Example 1.3 holds. Considering Example 1.3, it is natural to ask where does the gap lies between $R$ and $d$. We show a collapse on equilibria
$R_{\mathrm{DF}}(T)=d(T)$ (1.5)
for all AND–OR tree $T$: The randomized complexity with respect to the depth-first algorithms equals that with respect to the directional algorithms (RDAs). Here, no shape constraints are imposed on the tree $T$. Any node of $T$ can have any number (${\geq}2$) of child nodes, and there can be a mixture of AND and OR nodes among the child nodes. (1.5) is a strict equality without Landau symbol $O$. Thus, we obtain a separation result on the equilibria
$R(T)<R_{\mathrm{DF}}(T)$ (1.6)
for a certain $T$: The gap lies between the general algorithms and the depth-first algorithms. In other words, our result implies that there exists a case where a depth-first algorithm cannot attain the randomized complexity $R$. This study thus provides a deeper understanding of the depth-first algorithm for AND–OR trees.
@@ 2 Definitions and Preliminaries
In this section, we will define some basic concepts and see some fundamental facts on AND–OR trees. See also [3, Chapter 12] and [15] for helpful examples.
@@ 3.1 Replacement of sub-algorithms
[14, Lemma 3.1] There exists an RDA on $T$ that attains both $d_{0}(T)$ and $d_{1}(T)$.